# The Astrometric Resoeccentric Degeneracy: Eccentric Single Planets Mimic 2:1 Resonant Planet Pairs in Astrometry

### v1

@@ 5 Conclusion
If single-planet fits are assumed by default, circular resonant pairs may be mistaken for single eccentric planets – biasing inferred eccentricity distributions, dynamical interpretations, and occurrence rates. This underscores the need for orbit-fitting approaches that can handle multi-modal astrometric likelihoods, including this resoeccentric degeneracy. Samplers capable of robustly exploring degenerate orbital solutions are essential to avoid collapsing 2:1 configurations into single eccentric solutions. A combined astrometry and RV approach appears particularly promising. Systems like Gaia-4b and Gaia-5b – where RV follow-up already exists – offer ideal real-world test beds for assessing when and how RV observations can break the 2:1-eccentric astrometric degeneracy.
Several natural extensions follow from this work. It will be important to map out the regions of parameter space in which mutual inclination or follow-up measurements reliably distinguish the two orbital architectures, and to quantify the signal-to-noise thresholds neede for Gaia to do the same. Population-level indicators, such as anisotropies in the fitted argument of periastron distribution, may provide additional diagnostic power. Injection-recovery experiments that include resonant multi-planet architectures – in addition to single-planet systems – will be key for unbiased occurrence-rate measurements.
@@
We are very grateful to David Kipping for comments and feedback on an earlier version of this manuscript. We thank the CCA Astronomical Data Group and the Rice Research Group for helpful and illuminating conversations. D.A.Y. and T.L. are supported by Flatiron Research Fellowships at the Flatiron Institute, a division of the Simons Foundation. M.R. acknowledges support from Heising-Simons grants #2021-2802 and #2023-4478, as well as National Geographic grant EC-115062R-24 and NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant No. 80NSSC24K0153). ChatGPT was utilized to improve wording at the sentence level and assist with coding inquires – last accessed in 2025 November

### v2

@@ V Conclusion
If single planet fits are assumed by default, circular resonant pairs may be mistaken for single eccentric planets – biasing inferred eccentricity distributions, dynamical interpretations, and occurrence rates. This underscores the need for orbit-fitting approaches that can handle multi-modal astrometric likelihoods, including this resoeccentric degeneracy. Samplers capable of robustly exploring degenerate orbital solutions are essential to avoid collapsing 2:1 configurations into single eccentric solutions. A combined astrometry and RV approach appears particularly promising. Systems like Gaia-4b and Gaia-5b – where RV follow-up already exists – offer ideal real-world test beds for assessing when and how RV observations can break the 2:1-eccentric astrometric degeneracy.
Several natural extensions follow from this work. It will be important to map out the regions of parameter space in which mutual inclination or follow-up measurements reliably distinguish the two orbital architectures, and to quantify the signal-to-noise thresholds neede for Gaia to do the same. Population-level indicators, such as anisotropies in the fitted argument of periastron distribution, may provide additional diagnostic power. Injection-recovery experiments that include resonant multi-planet architectures – in addition to single planet systems – will be key for unbiased occurrence-rate measurements.
@@
We thank the anonymous reviewer who’s feedback improved this manuscript. We are very grateful to David Kipping for comments and feedback on an earlier version of this manuscript. We thank the CCA Astronomical Data Group and the Rice Research Group at Yale University for helpful and illuminating conversations. D.A.Y. is supported by a Juan Carlos Torres Postdoctoral Fellowship at the Massachusetts Institute of Technology. D.A.Y. and T.L. are supported by Flatiron Research Fellowships at the Flatiron Institute, a division of the Simons Foundation. M.R. acknowledges support from Heising-Simons grants #2021-2802 and #2023-4478, as well as National Geographic grant EC-115062R-24 and NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant No. 80NSSC24K0153). ChatGPT was utilized to improve wording at the sentence level and assist with coding inquires – last accessed in 2026 January.